# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
Note $U(\sigma)=\int A(\sigma(x))d^{3}x<\infty$ implies $A(2\sigma)\in L^{1}(\mathbb{R}^{3})$, and $\Omega_{>0}$ is bounded implies $A(2\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. Therefore, we know $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
We can then split $\mathbb{R}^{3}$ into $\mathbb{R}^{3}\setminus\Omega_{>0}$ and $\Omega_{>0}$, and show $\int_{\mathbb{R}^{3}}A\left(\sigma_{s}\right)\rightarrow\int_{\mathbb{R}^{3}}A(\sigma)$ as $s\rightarrow\infty$ by the monotone convergence theorem and dominated convergence theorem. Since $\Omega$ is bounded, by Hölder’s inequality we know $\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq C\|\tau\|_{L^{\infty}(\Omega)}$, then $\left\|\sigma_{s}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\left\|\sigma_{s}^{[1]}\right\|_{L^{\frac{4}{3}}(\Omega)}+\left\|\sigma_{s}^{[2]}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq$ $\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+C\|\tau\|_{L^{\infty}(\Omega)}$. It means $\left\{\sigma_{s}\right\}$ is a bounded set in $L^{\frac{4}{3}}(\Omega)$. Since $L^{\frac{4}{3}}(\Omega)$ is reflexive, thanks to [8, Theorem 3.18], there is a sequence $\left\{\sigma_{s_{n}}\right\}$ which converges weakly to a weak limit $\widetilde{\sigma}$. Since we know $\sigma_{s_{n}}\rightarrow\sigma$ a.e., by [55], we have $\widetilde{\sigma}=\sigma$. That is, $\sigma_{s_{n}}\rightharpoonup\sigma$ in $L^{\frac{4}{3}}(\Omega)$, therefore, we can apply what we did in the proof of Lemma 4.3, and obtain $G\left(\sigma_{s_{n}},\sigma_{s_{n}}\right)\rightarrow G(\sigma,\sigma),T_{J}\left(\sigma_{s_{n}}\right)\rightarrow T_{J}(\sigma)$. Collect the results above, we have $E_{J}(\sigma)=\lim\limits_{n\rightarrow\infty}E_{J}\left(\sigma_{s_{n}}\right)$. By the same arguments, we know for any sequence of $\left\{\sigma_{s_{m}}\right\}$, we can always pick a subsequence $\left\{\sigma_{s_{m_{k}}}\right\}$ which also satisfies $E_{J}(\sigma)=\lim\limits_{k\rightarrow\infty}E_{J}\left(\sigma_{s_{m_{k}}}\right)$. Thus, we get, $E_{J}(\sigma)=\lim\limits_{s\rightarrow\infty}E_{J}\left(\sigma_{s}\right)$ (one can show it by contradiction argument).
@@
In the proof of Theorem 4.14, the verification that $A(\sigma+|\tau|{L^{\infty}})\in L^{1}(\Omega{>0})$ relies only on the convexity of $A$, not on its specific form. Consequently, the same argument is applicable to more general pressure. For the polytropic case studied in this paper, one may also use the following estimate: $A(\sigma+\|\tau\|_{L^{\infty}})=\frac{K}{\gamma-1}(\sigma+\|\tau\|_{L^{\infty}})^{\gamma}\leq\frac{K}{\gamma-1}(2\max\{\sigma,\|\tau\|_{L^{\infty}}\})^{\gamma}=\frac{K}{\gamma-1}(2^{\gamma}\max\{\sigma^{\gamma},\|\tau\|_{L^{\infty}}^{\gamma}\})\leq\frac{K2^{\gamma}}{\gamma-1}(\sigma^{\gamma}+\|\tau\|_{L^{\infty}}^{\gamma})=2^{\gamma}(A(\sigma)+A(\|\tau\|_{L^{\infty}}))$.
@@
Remark 4.17.
Note that the continuity of $\rho$ has so far been established only on $\Omega$, not on the entire $\mathbb{R}^{3}$. This would suggest that there might be a sense of discontinuous density transition around the boundary for stars or planets. However, by additionally showing that the support lies in the interior, we will rule out this scenario and further extend the result to continuity on $\mathbb{R}^{3}$. Moreover, we have established the existence of a constrained minimizer, one can verify that such a minimizer satisfies the Euler–Lagrange equations (EL) on $\Omega$. However, to conclude that the reduced Euler–Poisson equations (EP’) hold in $\mathbb{R}^{3}$, we need to show the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$, particularly on $\partial\{\rho>0\}$. It turns out $W^{\infty}$ local energy minimizers can do this job as we mention in Remark 3.4, and we have Theorem 3.1. Thanks to Remark 2.11, it suffices to show that the minimizer can be placed inside the interior of $\Omega$, which will be discussed in the following sections.
@@ 5 Bound of the Supports of Density Functions
In this section we estimate a uniform bound for the supports of the density functions, under small mass ratio assumption which is desirable physically in star-planet systems. The analysis focuses on the planet’s density. For stars densities, the estimates are similar or even simpler, and can be adapted directly from the arguments in [46] and does not require the scaling method in subsection 5.1 and subsection 5.2. With these uniform bounds we will show in Section 6 that the minimizers can be located in the interior of the domains we define.
@@
$={\int{\frac{K{\left({\rho_{1-m}(x)+\rho_{m}(x)}\right)}^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(m)(x)\cdot\rho(m)(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+\frac{J^{2}}{2\left({I\left({\rho_{1-m}+\rho_{m}}\right)}\right)}$
@@
Notice since $\rho_{1-m}\in\Omega_{1-m}$, $\rho_{m}\in\Omega_{m}$, which distance is at least $\frac{\eta}{2}=\frac{J^{2}}{2m^{2}(1-m)^{2}}$ (24), we obtain $G\left({\rho_{1-m},\rho_{m}}\right)\leq\frac{2m\left({1-m}\right)}{\eta}<Cm^{3}$ for some $C>0$.
@@
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $I\left(\rho_{1-m}+\rho_{m}\right)>Cm^{-3}$.
@@
The proof is similar as in the proof of Lemma 4.9. ∎
@@
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $T_{J}(\rho(m))<Cm^{3}$.
@@
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $v(x,m)<Cm$.
@@
The proof is similar as in the proof of Lemma 4.10. ∎
@@
By the scaling method and uniqueness of the minimizer, we know $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ (see Theorem 3.11), where $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$ are the same as before, $\sigma$ is the non-rotating minimizer with mass 1. Similarly, $\sigma_{1-m}=\frac{1}{\widetilde{A}}\sigma\left(\frac{1}{\widetilde{B}}x\right)$, where $\widetilde{A}=(1-m)^{-\frac{2}{3\gamma-4}}$, $\widetilde{B}=(1-m)^{\frac{\gamma-2}{3\gamma-4}}$. Moreover, $\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\varrho_{m}\right)$, $\frac{{\widetilde{B}}^{3}}{{\widetilde{A}}^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{1-m}\right)=E_{0}\left(\varrho_{1-m}\right)$.
@@
Again, by Theorem 3.6, $\varrho_{1-m}$ and $\varrho_{m}$ have some nice properties, in particular they are spherically symmetry and radially decreasing after translations. Therefore, by comparison with the point masses, we get $G\left({\varrho_{1-m},\varrho_{m}}\right)=\mu\eta^{-1}=\frac{\mu^{3}}{J^{2}}$ (one can verify this by using Legendre polynomials or refer to [16, Corollary 2.31]), and $T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)\leq\frac{\mu^{3}}{2J^{2}}$ similarly as in Corollary 5.4. Thus, the right side of (40) is less than 0. This observation, together with Remark 5.7 gives us that:
@@
Therefore, we get $\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)<G\left({\rho_{1-m},\rho_{m}}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$, or $E_{0}\left(\widetilde{\rho_{m}}\right)<\frac{A^{\gamma}}{B^{3}}G\left({\rho_{1-m},\rho_{m}}\right)+E_{0}(\sigma)$. Since $\frac{A^{\gamma}}{B^{3}}=m^{\frac{6-5\gamma}{3\gamma-4}}$, $G\left({\rho_{1-m},\rho_{m}}\right)\leq Cm^{3}$ due to Remark 5.2, we have
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A\rho_{m}\left({Bx}\right)$ as in Definition 5.1, we have, up to translation, $\{\widetilde{\rho_{m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$. Here $\sigma$ is the non-rotating minimizer with mass 1.
@@
Due to Proposition 5.8, we know ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{m}}\right)}}=e_{0}<0$. Moreover, we can apply the concentration-compactness lemma of Lions [42, Theorem II.2 and Corollary II.1] to get that $\{\widetilde{\rho_{m}}\}$ is relatively compact up to translations in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. Thus, there is a subsequence $\{\widetilde{\rho_{m_{k}}}\}$ that converges, up to translations, to some function $f$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, $L^{1}\left(\mathbb{R}^{3}\right)$ and almost everywhere in $\mathbb{R}^{3}$. We denote again the subsequence after translations by $\left\{\widetilde{\rho_{m_{k}}}\right\}$. In particular, $\left\|f\right\|_{L^{1}}=1$.
@@
And we claim that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}$. In fact, we recall $E_{0}(\rho)={\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}$. For the inertial energy $\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}$, by $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. a.e. and Fatou’s Lemma we have ${\int{\frac{K\left({f(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}\leq\lim\limits_{k\rightarrow\infty}\inf{\int{\frac{K\left({\widetilde{\rho_{m_{k}}}(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}$. For the potential energy $\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}=\frac{1}{2}{\int{\rho(x)V_{\rho}(x)d^{3}x}}$, we have
@@
$\begin{split}&\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &=\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}+{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}}\right|+\left|{{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left\|{\widetilde{\rho_{m_{k}}}-f}\right\|_{L^{\frac{4}{3}}}\cdot\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}+\left\|f\right\|_{L^{\frac{12}{11}}}\cdot\left\|{V_{f}-V_{\widetilde{\rho_{m_{k}}}}}\right\|_{L^{12}}\end{split}$
@@
The last inequality above comes from the Hölder’s inequality. Thanks to Hardy-Littlewood-Sobolev Inequality Proposition A.3, we have $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$, and $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies $\left.V_{\widetilde{\rho_{m_{k}}}}\rightarrow V_{f}(x)\right$. in $L^{12}\left(\mathbb{R}^{3}\right)$. Thanks to the Interpolation Inequality, we have $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}\leq\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{1}}\cdot\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{4}{3}}}$, $\left\|f\right\|_{L^{\frac{12}{11}}}\leq\left\|f\right\|_{L^{1}}\cdot\left\|f\right\|_{L^{\frac{4}{3}}}$. Thus $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$ and $\left\|f\right\|_{L^{\frac{12}{11}}}$ are bounded, which—together with the inequality $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$—implies the boundedness of $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}$. Then
$\left.\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\rightarrow 0\right$.

### v2

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
Note $U(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma(x))\,dx<\infty$ implies $A(2\sigma)\in L^{1}(\mathbb{R}^{3})$, and $\Omega_{>0}$ is bounded implies $A(2\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. Therefore, we know $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
We can then split $\mathbb{R}^{3}$ into $\mathbb{R}^{3}\setminus\Omega_{>0}$ and $\Omega_{>0}$, and show $\int_{\mathbb{R}^{3}}A\left(\sigma_{s}\right)\,dx\rightarrow\int_{\mathbb{R}^{3}}A(\sigma)\,dx$ as $s\rightarrow\infty$ by the monotone convergence theorem and dominated convergence theorem. Since $\Omega$ is bounded, by Hölder’s inequality we know $\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq C\|\tau\|_{L^{\infty}(\Omega)}$, then
$\left\|\sigma_{s}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\left\|\sigma_{s}^{[1]}\right\|_{L^{\frac{4}{3}}(\Omega)}+\left\|\sigma_{s}^{[2]}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+C\|\tau\|_{L^{\infty}(\Omega)}$
It means $\left\{\sigma_{s}\right\}$ is a bounded set in $L^{\frac{4}{3}}(\Omega)$. Since $L^{\frac{4}{3}}(\Omega)$ is reflexive, thanks to [6, Theorem 3.18], there is a sequence $\left\{\sigma_{s_{n}}\right\}$ which converges weakly to a weak limit $\widetilde{\sigma}$. Since we know $\sigma_{s_{n}}\rightarrow\sigma$ a.e., by [31], we have $\widetilde{\sigma}=\sigma$. That is, $\sigma_{s_{n}}\rightharpoonup\sigma$ in $L^{\frac{4}{3}}(\Omega)$, therefore, we can apply what we did in the proof of Lemma 4.3, and obtain $G\left(\sigma_{s_{n}},\sigma_{s_{n}}\right)\rightarrow G(\sigma,\sigma)$, $T_{J}\left(\sigma_{s_{n}}\right)\rightarrow T_{J}(\sigma)$. Collecting the results above, we have $E_{J}(\sigma)=\lim\limits_{n\rightarrow\infty}E_{J}\left(\sigma_{s_{n}}\right)$. By the same arguments, we know for any sequence of $\left\{\sigma_{s_{m}}\right\}$, we can always pick a subsequence $\left\{\sigma_{s_{m_{k}}}\right\}$ which also satisfies $E_{J}(\sigma)=\lim\limits_{k\rightarrow\infty}E_{J}\left(\sigma_{s_{m_{k}}}\right)$. Thus, we get, $E_{J}(\sigma)=\lim\limits_{s\rightarrow\infty}E_{J}\left(\sigma_{s}\right)$ (one can show it by contradiction argument).
@@
In the proof of Theorem 4.14, the verification that $A(\sigma+|\tau|{L^{\infty}})\in L^{1}(\Omega{>0})$ relies only on the convexity of $A$, not on its specific form. Consequently, the same argument is applicable to more general pressure. For the polytropic case studied in this paper, one may also use the following estimate:
$A(\sigma+\|\tau\|_{L^{\infty}})$ $=\frac{K}{\gamma-1}(\sigma+\|\tau\|_{L^{\infty}})^{\gamma}$
$\leq\frac{K}{\gamma-1}(2\max\{\sigma,\|\tau\|_{L^{\infty}}\})^{\gamma}$
$=\frac{K2^{\gamma}}{\gamma-1}\max\{\sigma^{\gamma},\|\tau\|_{L^{\infty}}^{\gamma}\}$
$\leq\frac{K2^{\gamma}}{\gamma-1}(\sigma^{\gamma}+\|\tau\|_{L^{\infty}}^{\gamma})$
$=2^{\gamma}A(\sigma)+2^{\gamma}A(\|\tau\|_{L^{\infty}})$
@@ 5 Bound of the Supports of Density Functions
In this section we estimate a uniform bound for the supports of the density functions, under small mass ratio assumption which is physically desirable in star-planet systems. The analysis focuses on the planet’s density. For the star’s density, the estimates are similar or even simpler, and can be adapted directly from the arguments in [24] and do not require the scaling method in subsection 5.1 and subsection 5.2. With these uniform bounds we will show in Section 6 that the minimizers can be located in the interior of the domains we define.
@@
$={\int_{\mathbb{R}^{3}}{\frac{K{\left({\rho_{1-m}(x)+\rho_{m}(x)}\right)}^{\gamma}}{\gamma-1}\,dx}}-\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(m)(x)\cdot\rho(m)(y)}{\left|{x-y}\right|}\,dx\,dy}}+\frac{J^{2}}{2\left({I\left({\rho_{1-m}+\rho_{m}}\right)}\right)}$
@@
Notice since $\rho_{1-m}\in\Omega_{1-m}$, $\rho_{m}\in\Omega_{m}$, whose distance is at least $\frac{\eta}{2}=\frac{J^{2}}{2m^{2}(1-m)^{2}}$ (25), we obtain $G\left({\rho_{1-m},\rho_{m}}\right)\leq\frac{2m\left({1-m}\right)}{\eta}<Cm^{3}$ for some $C>0$.
@@
Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $I\left(\rho_{1-m}+\rho_{m}\right)>Cm^{-3}$.
@@
The proof is similar to the proof of Lemma 4.9. ∎
@@
Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $T_{J}(\rho(m))<Cm^{3}$.
@@
Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $v(x,m)<Cm$.
@@
The proof is similar to the proof of Lemma 4.10. ∎
@@
By the scaling method and uniqueness of the minimizer, we know $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ (see Theorem 3.9), where $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$ are the same as before, $\sigma$ is the non-rotating minimizer with mass 1. Similarly, $\sigma_{1-m}=\frac{1}{\widetilde{A}}\sigma\left(\frac{1}{\widetilde{B}}x\right)$, where $\widetilde{A}=(1-m)^{-\frac{2}{3\gamma-4}}$, $\widetilde{B}=(1-m)^{\frac{\gamma-2}{3\gamma-4}}$. Moreover, $\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\varrho_{m}\right)$, and $\frac{{\widetilde{B}}^{3}}{{\widetilde{A}}^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{1-m}\right)=E_{0}\left(\varrho_{1-m}\right)$.
@@
Again, by Theorem 3.6, $\varrho_{1-m}$ and $\varrho_{m}$ have some nice properties, in particular they are spherically symmetric and radially decreasing after translations. Therefore, by comparison with the point masses, we get $G\left({\varrho_{1-m},\varrho_{m}}\right)=\mu_{r}\eta^{-1}=\frac{\mu_{r}^{3}}{J^{2}}$ (one can verify this by using Legendre polynomials or refer to [24] [11, Section 2]), and $T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)\leq\frac{\mu_{r}^{3}}{2J^{2}}$ similar to that in Corollary 5.4. Thus, the right side of (41) is less than 0. This observation, together with Remark 5.7 gives us that:
@@
Therefore, we get $\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)<G\left({\rho_{1-m},\rho_{m}}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$ (i.e. $E_{0}\left(\widetilde{\rho_{m}}\right)<\frac{A^{\gamma}}{B^{3}}G\left({\rho_{1-m},\rho_{m}}\right)+E_{0}(\sigma)$). Since $\frac{A^{\gamma}}{B^{3}}=m^{\frac{6-5\gamma}{3\gamma-4}}$, $G\left({\rho_{1-m},\rho_{m}}\right)\leq Cm^{3}$ due to Remark 5.2, we have
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A\rho_{m}\left({Bx}\right)$ as in Definition 5.1, we have, up to translation, $\{\widetilde{\rho_{m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$. Here $\sigma$ is the non-rotating minimizer with mass 1.
@@
Due to Proposition 5.8, we know ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{m}}\right)}}=e_{0}<0$. Moreover, we can apply the concentration-compactness lemma of Lions [22, Theorem II.2 and Corollary II.1] to get that $\{\widetilde{\rho_{m}}\}$ is relatively compact up to translations in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. Thus, there is a subsequence $\{\widetilde{\rho_{m_{k}}}\}$ that converges, up to translations, to some function $f$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, $L^{1}\left(\mathbb{R}^{3}\right)$ and almost everywhere in $\mathbb{R}^{3}$. We denote again the subsequence after translations by $\left\{\widetilde{\rho_{m_{k}}}\right\}$. In particular, $\left\|f\right\|_{L^{1}}=1$.
@@
And we claim that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}$. In fact, we recall
$E_{0}(\rho)={\int_{\mathbb{R}^{3}}{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}\,dx}}-\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}\,dx\,dy}}$
For the inertial energy $\int_{\mathbb{R}^{3}}{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}\,dx}$, by $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. a.e. and Fatou’s Lemma we have ${\int_{\mathbb{R}^{3}}{\frac{K\left({f(x)}\right)^{\gamma}}{\gamma-1}\,dx}}\leq\lim\limits_{k\rightarrow\infty}\inf{\int_{\mathbb{R}^{3}}{\frac{K\left({\widetilde{\rho_{m_{k}}}(x)}\right)^{\gamma}}{\gamma-1}\,dx}}$.
For the potential energy $\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}\,dx\,dy}}=\frac{1}{2}{\int_{\mathbb{R}^{3}}{\rho(x)V_{\rho}(x)\,dx}}$, we have
@@
$\begin{split}&\quad\,\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &=\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}+{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &\leq\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}}\right|+\left|{{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &\leq\left\|{\widetilde{\rho_{m_{k}}}-f}\right\|_{L^{\frac{4}{3}}}\cdot\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}+\left\|f\right\|_{L^{\frac{12}{11}}}\cdot\left\|{V_{f}-V_{\widetilde{\rho_{m_{k}}}}}\right\|_{L^{12}}\end{split}$
@@
The last inequality above comes from Hölder’s inequality. Thanks to Hardy-Littlewood-Sobolev Inequality (Proposition A.1), we have $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$, and $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies $\left.V_{\widetilde{\rho_{m_{k}}}}\rightarrow V_{f}(x)\right$. in $L^{12}\left(\mathbb{R}^{3}\right)$. Due to the Interpolation Inequality [6, Section 4.2], we have $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}\leq\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{1}}\cdot\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{4}{3}}}$, and $\left\|f\right\|_{L^{\frac{12}{11}}}\leq\left\|f\right\|_{L^{1}}\cdot\left\|f\right\|_{L^{\frac{4}{3}}}$. Thus $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$ and $\left\|f\right\|_{L^{\frac{12}{11}}}$ are bounded, which—together with the inequality $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$—implies the boundedness of $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}$. Then
$\left.\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\rightarrow 0\right$.